OpenStudy (anonymous):

What are the diameter and radius of a circle with a circumference of 28π m? A. d = 28 m r = 14 m B. d = 14 m r = 28 m C. d = 14 m r = 7 m D. d = 7 m r = 14 m

OpenStudy (anonymous):

d=c/π d=28πm/π d=28m r=d/2 r=28m/2 r=14

OpenStudy (anonymous):

I mean 14 m

OpenStudy (anonymous):

Which is the best approximation for the circumference of a circular fish pond that is 24 ft in diameter? A. 38 ft B. 72 ft C. 75 ft D. 150 ft

OpenStudy (anonymous):

rdy?

OpenStudy (anonymous):

c=d•π c=24•π c≈75.398≈75

OpenStudy (anonymous):

What is the radius of of a circle with a circumference of 18π ft? A. 6 ft B. 9 ft C. 18 ft D. 9π ft

OpenStudy (anonymous):

d=c/π d=18π/π d=18 r=d/2 r=18/2 r=9